# The Sommerfeld-Rellich Framework for Scattering on Hyperbolic Space: Far-Field Patterns and Inverse Problems

### v2

@@ Abstract.
This paper establishes a complete time-harmonic scattering theory for hyperbolic space, formulating it within the classical Sommerfeld-Rellich paradigm centered on far-field patterns–a foundational framework that has been absent despite the well-developed spectral and time-dependent theories for this geometry. We explicitly construct the ingoing and outgoing fundamental solutions for the Helmholtz operator and perform a precise asymptotic analysis at the conformal boundary to derive a hyperbolic Sommerfeld radiation condition. This condition, which is a local criterion at infinity, uniquely selects physically admissible outgoing solutions. We prove a fhyperbolic Rellich theorem guaranteeing the uniqueness of the scattered field and its far-field pattern from asymptotic data. Within this rigorous framework, we solve the direct scattering problem for compact sources, penetrable media, and impenetrable obstacles, providing explicit representations for the corresponding far-field patterns. As a principal application and demonstration of the framework’s utility, we initiate the study of inverse scattering on hyperbolic space, formulating both the inverse obstacle and inverse medium problems where the objective is to reconstruct the scatterer from measurements of its far-field pattern. Our work provides the essential theoretical underpinning for a far-field-based approach to scattering and inversion in hyperbolic geometry and lays the groundwork for extensions to asymptotically hyperbolic manifolds.
@@
Mathematics Subject Classification. Primary: 35P25, 58J50; Secondary: 35R30, 43A85
The first author was partly supported by the National Key Research and Development Program (No. 2022YFA1006900) and National Natural Science Foundation of China (No. 12271027). The second author was supported by ANR/RGC Joint Research Scheme, A-CityU203/19, and the Hong Kong RGC General Research Funds (projects (projects 11311122, 11304224 and 113008).
@@ 1.1. Background and motivation
This body of work has largely been developed within the time-dependent or stationary spectral framework, where the primary objects are wave operators, the scattering matrix, or the boundary resolvent. In contrast, the time-harmonic framework—centered on the explicit characterization of outgoing radiation conditions, far-field patterns, and their role as measurable data for inverse problems—has been the cornerstone of Euclidean scattering theory since the seminal work of Sommerfeld [37] and Rellich [36]. Sommerfeld’s radiation condition (1912) and Rellich’s uniqueness theorem (1943) provided the rigorous foundation for the exterior Helmholtz equation. The systematic use of the far-field pattern as the central data for inverse scattering was later developed by Colton, Kress, and others [5], creating a complete “Sommerfeld-Rellich-Colton-Kress” (SRCK) structure. This structure is not merely theoretical; it is the operational foundation of most numerical reconstruction algorithms in acoustics, electromagnetics, and medical imaging, where multi-static far-field data serves as the primary input.
@@ 2. The time-harmonic scattering theory in Euclidean space
In this section, we briefly introduce the classical time‑harmonic scattering framework in Euclidean space, governed by the Helmholtz equation. Our purpose is twofold: to fix notation and conventions, and more importantly, to recall the key structural elements—radiation conditions, far‑field patterns, well‑posedness, and uniqueness results—that will guide our construction of an analogous theory on hyperbolic space. For comprehensive treatments of Euclidean scattering we refer to the monographs [4, 5]; the time‑dependent theory is discussed, for example, in [10, 23, 24].
@@ 3.1.2. Poincaré ball model
The advantage of Poincare ball model lies in clear expression of differential operator and is used widely in the fields of analysis and PDEs. The establishment of our time-harmonic scattering theory for hyperbolic space mainly use the Poincare ball model.
@@ 3.1.3. Half-space model
The advantage half space model lies in clear scaling transform and sharp geometric inequality such as Hardy-Sobolv-Mayza inequality on the half space can be equivalently seen as Poincare-Sobolev inequality on the hyperbolic space.
@@ 3.4. Green’s formulas on bounded hyperbolic domains
Let $B_{\mathbb{B}^{n}}(0,R)$ be the hyperbolic ball of radius $R$ centered at 0. For $u,v\in C^{2}(\overline{B_{\mathbb{B}^{n}}(0,R)})$,
@@
Recall the conformal relation between the hyperbolic Laplacian and the Euclidean Laplacian:
@@ 4.2.1. Analytic continuation from the shifted Laplacian
Observe that the Helmholtz operator $\mathcal{L}_{\mu}$ corresponds formally to taking $z=\pm\mathrm{i}\mu$, which lies outside the original domain of $z^{2}\in\mathbb{C}\setminus(-\infty,0)$ and $Re(z)>-\frac{1}{2}$. However, the analytically continued family
@@ 4.4. Radiation condition and far-field pattern
as $\rho\rightarrow+\infty$. We called the condition (4.10) hyperbolic radiation condition and prove that this radiation condition will make sure the existence and uniqueness for solutions of the wave equation involving the source or potential later.
@@
4.5. Scattering theory for Helmholtz equation with the source
@@
$f(y)=D_{n}\left(\frac{A_{n,-\frac{\mu\mathrm{i}}{2}}}{2^{\frac{n-1}{2}}}\right)^{-1}\left(\int_{0}^{\pi}(\sin t)^{-\mu\mathrm{i}}dt\right)^{-1}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}u_{\infty}^{\mu/2}(\hat{x})e_{\mu,\hat{x}}(y)|c(\mu)|^{-2}d\mu d\sigma.$
That is to say that if we know the information of far-field pattern $u_{\infty}(\hat{x},\frac{\mu}{2})$ for all observable direction $\hat{x}$ and frequency $\mu$, then the source $f(x)$ can be uniquely determined by inverse Fourier transform in hyperbolic space. This is one of the rare cases that unknown source can be expressed explicitly.
@@
(4.15) $\begin{split}&\lim\limits_{R\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}|\frac{\partial u}{\partial\rho}-\left(\mu i-\frac{n-1}{2}\right)\tanh(\frac{\rho}{2})u|^{2}d\sigma_{\mathbb{B}^{n}}\\ &\ \ =\lim\limits_{R\rightarrow+\infty}\int_{\partial B^{n}(0,\tanh(\frac{R}{2}))}\left|\frac{\partial u}{\partial\rho}-\left(\mu i-\frac{n-1}{2}\right)\tanh(\frac{\rho}{2})u\right|^{2}\left(\frac{2}{1-|x|^{2}}\right)^{n-1}d\sigma\\ &\ \ =\lim\limits_{R\rightarrow+\infty}\left(2\cosh(\frac{R}{2})\right)^{n-1}\int_{\partial B^{n}(0,\tanh(\frac{R}{2}))}o\left(\frac{1}{\sinh^{n-1}(\rho)}\right)d\sigma=0.\end{split}$
@@ 4.7. Scattering theory for Helmholtz equation involving potential
If imaginary part $\mathbb{\Im}(q)$of the potential $q$ is non-negative, then $u$ is existing and unique.
@@
We have proved that $u$ is existing and unique. Furthermore, the total wave $u$ can be written as far-field pattern form. In fact, for any $x\in\mathbb{B}^{n}$, applying Green representative formula and the accurate expansion of $G_{-\mu i}$, we have
@@ 4.8. Scattering theory for obstacle problems in hyperbolic space
(4.46) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u-\frac{(n-1)^{2}}{4}u=\mu^{2}u,\ \ x\in\mathbb{B}^{n}\setminus\Omega\\ &\mathcal{B}(u)=0,\ \ x\in\partial\Omega,\end{cases}$
where $\mathcal{B}$ denotes the boundary operator, $u(x)=u^{i}+u^{s}(x)$, $u^{i}(x)=e_{-2\mu,\xi}(x)$ is an incident wave satisfying
$\Delta_{\mathbb{B}^{n}}u^{i}-\frac{(n-1)^{2}}{4}u^{i}=\mu^{2}u^{i}$
@@
$\frac{\partial u^{s}}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u^{s}=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).$
@@
(4.47) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u^{s}-\frac{(n-1)^{2}}{4}u^{s}=\mu^{2}u^{s},\ \ x\in\mathbb{B}^{n}\setminus\Omega.\\ &\mathcal{B}(u^{s}+u^{i})=0,\ \ x\in\partial\Omega\\ &\frac{\partial u_{s}}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u_{s}=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).\end{cases}$
@@
The total wave $u$ is existing and unique.
@@
(4.48) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u-\frac{(n-1)^{2}}{4}u=\mu^{2}u,\ \ x\in\mathbb{B}^{n}\setminus\Omega\\ &u=0,\ \ x\in\partial\Omega,\\ &\frac{\partial(u-u^{i})}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})(u-u^{i})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).\end{cases}$
@@
(4.49) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}(u_{1}-u_{2})-\frac{(n-1)^{2}}{4}(u_{1}-u_{2})=\mu^{2}(u_{1}-u_{2}),\ \ x\in\mathbb{B}^{n}\setminus\Omega.\\ &u_{1}-u_{2}=0,\ \ x\in\partial\Omega,\\ &\frac{\partial(u_{1}-u_{2})}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})(u_{1}-u_{2})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).\end{cases}$
@@
Applying the hyperbolic Rellich theorem, we derive that $u_{1}-u_{2}$ is equal to zero in $\mathbb{B}^{n}\setminus\Omega$. This accomplishes the proof of uniqueness. Now, we start to show the existence. Since $u(x)=u^{s}(x)+e_{-2\mu,\xi}(x)$, we only need to show that $u^{s}$ is existing. We need the following lemma.
@@
(4.54) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u^{s}-\frac{(n-1)^{2}}{4}u^{s}=\mu^{2}u^{s},\ \ x\in\mathbb{B}^{n}\setminus\Omega,\\ &u^{s}=-e_{2\mu,\xi}(x),\ \ x\in\partial\Omega,\\ &\frac{\partial u^{s}}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u^{s}=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right),\end{cases}$
@@
(4.55) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u_{1}-\frac{(n-1)^{2}}{4}u_{1}=\mu^{2}u_{1},\ \ x\in B_{\mathbb{B}^{n}}(0,R_{0})\setminus\Omega,\\ &u_{1}=-e_{2\mu,\xi}(x),\ \ x\in\partial\Omega,\\ &\frac{\partial u_{1}}{\partial\nu_{\mathbb{B}^{n}}}=\Lambda u_{1},\ \ x\in\partial B_{\mathbb{B}^{n}}(0,R_{0}),\end{cases}$
@@
(4.56) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}W-\frac{(n-1)^{2}}{4}W=\mu^{2}W,\ \ x\in\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R_{0}),\\ &W=\psi\in H^{\frac{1}{2}}(\partial B_{\mathbb{B}^{n}}(0,R_{0})),\\ &\frac{\partial W}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})W=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).\end{cases}$
@@
(4.57) $\begin{split}u_{1}(x)&=-\int_{\partial\Omega}\left(\frac{\partial u_{1}(y)}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\Lambda u_{1}(y)G_{-\mu\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)\end{split}$
@@
$\int_{\partial B_{\mathbb{B}^{n}}(0,R_{0})}\left(\Lambda u_{1}(y)G_{-\mu\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)=0.$
For $x\in B_{\mathbb{B}^{n}}(0,R_{0})$, since both $G_{-\mu\mathrm{i}}(\rho(x,.))$ and $W(.)$ satisfy equation
$\Delta_{\mathbb{B}^{n}}u-\frac{(n-1)^{2}}{4}u=\mu^{2}u,\ \ y\in\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R_{0}).$
@@
Since $W$ and $G_{-\mu\mathrm{i}}$ satisfy hyperbolic radiation condition, then we deduce that
$\lim\limits_{R\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left|\left(\frac{\partial W}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)W\right|^{2}d\sigma_{\mathbb{B}^{n}}=0$
@@
$\lim\limits_{R\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left|\left(\frac{\partial G_{-\mu\mathrm{i}}}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)G_{-\mu\mathrm{i}}\right|^{2}d\sigma_{\mathbb{B}^{n}}=0.$
@@
$\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\Lambda u_{1}(y)G_{-\mu\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)=0.$

### v3

@@ Abstract.
This paper establishes a complete time-harmonic scattering theory for hyperbolic space, formulating it within the classical Sommerfeld-Rellich paradigm centered on far-field patterns–a foundational framework that has been absent despite the well-developed spectral and time-dependent theories for this geometry. We explicitly construct the ingoing and outgoing fundamental solutions for the Helmholtz operator and perform a precise asymptotic analysis at the conformal boundary to derive a hyperbolic Sommerfeld radiation condition. This condition, which is a local criterion at infinity, uniquely selects physically admissible outgoing solutions. We prove a hyperbolic Rellich theorem guaranteeing the uniqueness of the scattered field and its far-field pattern from asymptotic data. Within this rigorous framework, we solve the direct scattering problem for compact sources, penetrable media, and impenetrable obstacles, providing explicit representations for the corresponding far-field patterns. As a principal application and demonstration of the framework’s utility, we initiate the study of inverse scattering on hyperbolic space, formulating both the inverse obstacle and inverse medium problems where the objective is to reconstruct the scatterer from measurements of its far-field pattern. Our work lays the groundwork for a far-field-based approach to scattering and inversion in hyperbolic geometry. Due to the local feature of the hyperbolic Sommerfeld radiation condition, our study can be readily extended to the broader asymptotically hyperbolic manifolds.
@@
Mathematics Subject Classification (2010): 35R30, 35L05.
The first author was partly supported by the National Key Research and Development Program (No. 2022YFA1006900) and National Natural Science Foundation of China (No. 12271027), the second author was supported by NSFC/RGC Joint Research Scheme, N_CityU101/21, ANR/RGC Joint Research Scheme, A-CityU203/19, and the Hong Kong RGC General Research Funds (projects 11311122, 12301420 and 11300821).
@@ 1.1. Background and motivation
This body of work has largely been developed within the time-dependent or stationary spectral framework, where the primary objects are wave operators, the scattering matrix, or the boundary resolvent. In contrast, the time-harmonic framework—centered on the explicit characterization of outgoing radiation conditions, far-field patterns, and their role as measurable data for inverse problems—has been the cornerstone of Euclidean scattering theory since the seminal work of Sommerfeld [34] and Rellich. Sommerfeld’s radiation condition (1912) and Rellich’s uniqueness theorem (1943) provided the rigorous foundation for the exterior Helmholtz equation. The systematic use of the far-field pattern as the central data for inverse scattering was later developed by Colton, Kress, and others [5], creating a complete “Sommerfeld-Rellich-Colton-Kress” (SRCK) structure. This structure is not merely theoretical; it is the operational foundation of most numerical reconstruction algorithms in acoustics, electromagnetics, and medical imaging, where multi-static far-field data serves as the primary input.
@@ 2. The time-harmonic scattering theory in Euclidean space
In this section, we provide a brief overview of the classical time-harmonic scattering framework in Euclidean space—or more generally, on noncompact manifolds with Euclidean ends (see the remark below)—governed by the Helmholtz equation. Our aim is twofold: to establish notation and conventions, and more importantly, to highlight the essential structural features—such as radiation conditions, far-field patterns, well-posedness, and uniqueness results—that will serve as a guide for constructing an analogous theory in hyperbolic space. For comprehensive treatments of Euclidean scattering we refer to the monographs [4, 5]; the time‑dependent theory is discussed, for example, in [10, 23].
@@ 2.1. Direct scattering problems
In the discussion above, we focused primarily on time-harmonic scattering from compactly supported inhomogeneities in Euclidean space. However, given that the Sommerfeld radiation condition (2.4) is formulated locally, the theory can be readily extended to noncompact manifolds with Euclidean ends (cf. [2]).
@@ 3.1.2. Poincaré ball model
The advantage of Poincare ball model lies in clear expression of differential operator and is used widely in the fields of analysis and PDEs. The establishment of our time-harmonic scattering theory for hyperbolic space mainly uses the Poincare ball model.
@@ 3.1.3. Half-space model
The advantage half space model lies in clear scaling and sharp geometric inequality such as Hardy-Sobolv-Mayza inequality on the half space can be equivalently seen as Poincare-Sobolev inequality on the hyperbolic space.
@@ 3.4. Green’s formulas on bounded hyperbolic domains
Let $B_{\mathbb{B}^{n}}(0,R)$ be the hyperbolic ball of radius $R$ centered at 0. For $u,v\in C^{2}(\overline{B_{\mathbb{B}^{n}}(0,R)})$, there holds
@@
Recall the conformal laws between the hyperbolic Laplacian and the Euclidean Laplacian:
@@ 4.2.1. Analytic continuation from the shifted Laplacian
Observe that the Helmholtz operator $\mathcal{L}_{\mu}$ corresponds formally to taking $z=\pm\mathrm{i}\mu$, which lies outside the original domain of definition $k>0$. However, the analytically continued family
@@ 4.4. Radiation condition and far-field pattern
as $\rho\rightarrow+\infty$. We called the condition (4.10) hyperbolic radiation condition and prove that this radiation condition will guarantee the existence and uniqueness for solutions of the wave equation involving the source or potential later.
@@
4.5. Scattering theory for Helmholtz equation with a source
@@
$f(y)=D_{n}\left(\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\right)^{-1}\left(\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\right)^{-1}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}u_{\infty}^{\mu/2}(\hat{x})e_{\mu,\hat{x}}(y)|c(\mu)|^{-2}d\mu d\sigma.$
That is to say that if we know the information for far-field pattern $u_{\infty}(\hat{x},\frac{\mu}{2})$ for all observable direction $\hat{x}$ and frequency $\mu$, then the source $f(x)$ can be uniquely determined by inverse Fourier transform in hyperbolic space. This is one of the rare cases that unknown source can be expressed explicitly.
@@
(4.15) $\begin{split}&\lim\limits_{R\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\bigg|\frac{\partial u}{\partial\rho}-\left(\mu i-\frac{n-1}{2}\right)\tanh(\frac{\rho}{2})u\bigg|^{2}d\sigma_{\mathbb{B}^{n}}\\ &\ \ =\lim\limits_{R\rightarrow+\infty}\int_{\partial B^{n}(0,\tanh(\frac{R}{2}))}\left|\frac{\partial u}{\partial\rho}-\left(\mu i-\frac{n-1}{2}\right)\tanh(\frac{\rho}{2})u\right|^{2}\left(\frac{2}{1-|x|^{2}}\right)^{n-1}d\sigma\\ &\ \ =\lim\limits_{R\rightarrow+\infty}\left(2\cosh(\frac{R}{2})\right)^{n-1}\int_{\partial B^{n}(0,\tanh(\frac{R}{2}))}o\left(\frac{1}{\sinh^{n-1}(\rho)}\right)d\sigma=0.\end{split}$
@@ 4.7. Scattering theory for Helmholtz equation involving potential
If imaginary part $\mathbb{\Im}(q)$of the potential $q$ is non-negative, then there exits a unique solution $u\in H^{1}_{loc}(\mathbb{B}^{n})$.
@@
We have proved that $u$ exists and is unique. Furthermore, the total wave $u$ can be written as far-field pattern form. In fact, for any $x\in\mathbb{B}^{n}$, applying Green’s representation formula and the accurate expansion of $G_{-\mu i}$, we have
@@ 4.8. Scattering theory for obstacle problems in hyperbolic space
(4.46) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u-\frac{(n-1)^{2}}{4}u=\mu_{0}^{2}u,\ \ x\in\mathbb{B}^{n}\setminus\Omega\\ &\mathcal{B}(u)=0,\ \ x\in\partial\Omega,\end{cases}$
where $\mathcal{B}$ denotes the boundary operator, $u(x)=u^{i}+u^{s}(x)$, $u^{i}(x)=e_{-2\mu_{0},\xi}(x)$ is an incident wave satisfying
$\Delta_{\mathbb{B}^{n}}u^{i}-\frac{(n-1)^{2}}{4}u^{i}=\mu_{0}^{2}u^{i}$
@@
$\frac{\partial u^{s}}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u^{s}=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right),$
@@
(4.47) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u^{s}-\frac{(n-1)^{2}}{4}u^{s}=\mu_{0}^{2}u^{s},\ \ x\in\mathbb{B}^{n}\setminus\Omega.\\ &\mathcal{B}(u^{s}+u^{i})=0,\ \ x\in\partial\Omega\\ &\frac{\partial u_{s}}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u_{s}=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).\end{cases}$
@@
There exists a unique solution $u\in H_{loc}^{1}(\mathbb{B}^{n})$ to the scattering system (4.47).
@@
(4.48) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u-\frac{(n-1)^{2}}{4}u=\mu_{0}^{2}u,\ \ x\in\mathbb{B}^{n}\setminus\Omega\\ &u=0,\ \ x\in\partial\Omega,\\ &\frac{\partial(u-u^{i})}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})(u-u^{i})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).\end{cases}$
@@
(4.49) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}(u_{1}-u_{2})-\frac{(n-1)^{2}}{4}(u_{1}-u_{2})=\mu_{0}^{2}(u_{1}-u_{2}),\ \ x\in\mathbb{B}^{n}\setminus\Omega.\\ &u_{1}-u_{2}=0,\ \ x\in\partial\Omega,\\ &\frac{\partial(u_{1}-u_{2})}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})(u_{1}-u_{2})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right).\end{cases}$
@@
Applying the hyperbolic Rellich theorem, we derive that $u_{1}-u_{2}$ is equal to zero in $\mathbb{B}^{n}\setminus\Omega$. This accomplishes the proof of uniqueness. Now, we start to show the existence. Since $u(x)=u^{s}(x)+e_{-2\mu,\xi}(x)$, we only need to show that the existence of the scattered wave $u^{s}$. We need the following lemma.
@@
(4.54) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u^{s}-\frac{(n-1)^{2}}{4}u^{s}=\mu_{0}^{2}u^{s},\ \ x\in\mathbb{B}^{n}\setminus\Omega.\\ &u^{s}=-e_{2\mu,\xi}(x),\ \ x\in\partial\Omega\\ &\frac{\partial u_{s}}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u_{s}=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right)\end{cases}$
@@
(4.55) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u_{1}-\frac{(n-1)^{2}}{4}u_{1}=\mu_{0}^{2}u_{1},\ \ x\in B_{\mathbb{B}^{n}}(0,R_{0})\setminus\Omega.\\ &u_{1}=-e_{2\mu,\xi}(x),\ \ x\in\partial\Omega\\ &\frac{\partial u_{1}}{\partial\nu_{\mathbb{B}^{n}}}=\Lambda u_{1},\ \ x\in\partial B_{\mathbb{B}^{n}}(0,R_{0})\end{cases}$
@@
$\begin{cases}&-\Delta_{\mathbb{B}^{n}}W-\frac{(n-1)^{2}}{4}W=\lambda_{0}^{2}W,\ \ x\in\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R_{0}).\\ &W=\psi\in H^{\frac{1}{2}}(\partial B_{\mathbb{B}^{n}}(0,R_{0}))\\ &\frac{\partial W}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})W=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right)\end{cases}$
@@
$\begin{split}u_{1}(x)&=-\int_{\partial\Omega}\left(\frac{\partial u_{1}(y)}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu_{0}\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu_{0}\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\Lambda u_{1}(y)G_{-\mu_{0}\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu_{0}\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)\end{split}$
@@
$\int_{\partial B_{\mathbb{B}^{n}}(0,R_{0})}\left(\Lambda u_{1}(y)G_{-\mu_{0}\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu_{0}\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)=0.$
For $x\in B_{\mathbb{B}^{n}}(0,R_{0})$, since both $G_{-\mu_{0}\mathrm{i}}(\rho(x,.))$ and $W(.)$ satisfy equation
$\Delta_{\mathbb{B}^{n}}u-\frac{(n-1)^{2}}{4}u=\mu_{0}^{2}u,\ \ y\in\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R_{0}).$
@@
Since $W$ and $G_{-\mu_{0}\mathrm{i}}$ satisfy hyperbolic radiation condition, then we deduce that
$\lim\limits_{R\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left|\left(\frac{\partial W}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)W\right|^{2}d\sigma_{\mathbb{B}^{n}}=0$
@@
$\lim\limits_{R\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left|\left(\frac{\partial G_{-\mu_{0}\mathrm{i}}}{\partial\rho}-(\mu_{0}\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)G_{-\mu_{0}\mathrm{i}}\right|^{2}d\sigma_{\mathbb{B}^{n}}=0.$
@@
$\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\Lambda u_{1}(y)G_{-\mu_{0}\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu_{0}\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y)=0.$